# Optimizing Multi-modality Trackers via Sensitivity-regularized Tuning

### v2

@@ Abstract
This paper tackles the critical challenge of optimizing multi-modality trackers by effectively adapting pre-trained models for RGB data. Existing fine-tuning paradigms oscillate between excessive freedom and over-restriction, both leading to a suboptimal plasticity-stability trade-off. To mitigate this dilemma, we propose a novel sensitivity-regularized fine-tuning framework, which delicately refines the learning process by incorporating intrinsic parameter sensitivities. Through a comprehensive investigation of the transition from pre-trained to multi-modal contexts, we identify that parameters sensitive to pivotal foundational patterns and cross-domain shifts are the primary drivers of this issue. Specifically, we first probe the tangent space of pre-trained weights to measure and orient prior sensitivities, dedicated to preserving generalization. Subsequently, we characterize transfer sensitivities during the tuning phase, emphasizing adaptability and stability. By incorporating these sensitivities as unified regularization terms, our method significantly enhances the transferability across modalities. Extensive experiments showcase the superior performance of our method, surpassing current state-of-the-art techniques across various multi-modality tracking benchmarks. The source code and models will be publicly available at https://github.com/zhiwen-xdu/SRTrack.
@@ I Introduction
Object tracking, a foundation task of visual perception, has seen significant advancements over the past decades [16, 54, 67, 4]. Despite the promising results, RGB-based trackers often struggle with some complex and degraded conditions, such as extreme illumination, motion blur, and occlusions. Therefore, multi-modality tracking with more comprehensive information (e.g., event, depth, thermal) has garnered growing interest. With the popularity of the data-driven methods in the object tracking community, both data scale and model size have experienced huge explosions [58, 29, 10, 9]. There is a prevailing paradigm that explores pre-trained trackers on large-scale RGB-based datasets and adapts them to diverse auxiliary modalities, a process known as cross-modal fine-tuning or transfer learning, to enhance performance and accelerate convergence.
@@
Some existing approaches have explored the full fine-tuning (FFT) paradigm [45, 73, 41], where trackers are initialized with pre-trained weights and subsequently optimized for task-specific objectives. Nevertheless, due to the significant domain gaps and limited scale of auxiliary modalities, these methods struggle to preserve the pre-trained knowledge structure during transfer, often inducing severe overfitting (i.e., a widened train-test gap and deteriorating test performance). In contrast to full fine-tuning, recent research has shifted toward parameter efficient fine-tuning (PEFT) [18, 21, 7]. The PEFT keeps the majority of pre-trained parameters frozen, updating only a small fraction of modality-specific ones to retain prior knowledge. Several methods fall under this umbrella [70, 17, 51], including prompt tuning, visual adapter, etc. Although effective, PEFT-based methods impose rigid constraints on the primary model weights, resulting in underfitting (i.e., a restricted upper bound on training performance) when handling the significant distribution drifts. In summary, existing fine-tuning paradigms fluctuate between excessive flexibility and over-restriction, both contributing to misfitting and a sub-optimal plasticity-stability trade-off between pre-trained knowledge and downstream adaptation (as Fig. 1).
@@
In this work, we endeavor to mitigate the misfitting dilemma in cross-modal tracker adaptation by strategically regularizing the learning process. By probing parameter responses spanning from pre-training to fine-tuning phases, we identify that pronounced parameter sensitivities are the primary cause of degraded prior generalizability and transfer adaptability. To mitigate this, we propose a sensitivity-regularized fine-tuning framework that delicately calibrates the gradient updates to boost model transfer with precision. Specifically, we optimize multi-modality trackers from the following perspectives.
  1. (1)
@@
    Formulating Prior Sensitivity. We commence by investigating the tangent space of pre-trained parameters, utilizing it as a critical indicator to assess and preserve prior generalizability for downstream fine-tuning. Furthermore, we estimate this sensitivity via an eigen-decomposition approximation.
  2. (2)
@@
    Modeling Transfer Sensitivity. To further elucidate the adaptation challenges, we explore how sparse gradients exacerbate tuning instability. Accordingly, we formulate transfer sensitivity using off-the-shelf gradient matrices, aiming to facilitate effective gradient rebalancing.
  3. (3)
@@
    Sensitivity Regularized Tuning. By harnessing the identified parameter sensitivities, we suggest an adaptive tuning scheme that safeguards essential pre-trained knowledge and fosters coherent multi-modality representations through finely modulated, sensitivity-driven updates. This mechanism facilitates seamless transfer across various multi-modality tracking tasks, continuously enhancing the model during the training phase.
@@
Our method strategically guides the cross-domain fine-tuning process to optimize downstream tracking tasks. Extensive experimental results showcase our method achieves new state-of-the-art results across three multi-modality tracking tasks (RGB-Event, RGB-Depth, RGB-Thermal) and seven benchmarks, spanning diverse pre-trained tracking models. Comprehensive ablation studies demonstrate the effectiveness of the sensitivity-aware regularization fine-tuning concept.
In summary, the main contributions of this paper are:
@@
  • we revisit the misfitting issue of multi-modality tracking for adapting foundation models and propose a novel regularized tuning framework to indicate better transferability, which is orthogonal to the existing FFT and PEFT methods;
@@
  • we formulate the parameter sensitivity with respect to pre-trained knowledge and transfer stability, and introduce a sensitivity-aware update strategy to refine the learning process, thereby facilitating the generalization and adaptability of trackers;
@@
The remainder of the paper is organized as follows. Section II reviews existing literature on multi-modal trackers and cross-modal transfer learning methods. In Section III, we rigorously formulate prior and transfer sensitivities from both pre-training and fine-tuning perspectives, followed by the implementation of sensitivity-regularized tuning. Section IV presents extensive experiments to demonstrate the effectiveness of our method, along with comprehensive ablation studies to analyze the impact of different components and designs. Finally, Section V concludes the paper.
II Related Work
II-A Multi-Modal Object Tracking
@@
Object tracking, a cornerstone task in computer vision, involves predicting the position and scale of an object in future frames based on its initial appearance [48, 2, 28, 68]. However, the unreliability of RGB-only paradigms to adverse conditions has catalyzed research into multi-modal tracking, where auxiliary signals (e.g., event, depth, thermal data) are integrated to remedy the intrinsic deficiencies of visible imagery [66]. For instance, event cameras provide robust dynamic scene information under extreme motion or lighting. Leveraging this, [72, 11, 46, 63, 43, 44] combined RGB and event streams to predict objects in high-speed and low-dynamic scenarios. Meanwhile, depth cues offer geometric priors that are particularly useful for handling occlusion and background clutter. [32, 36, 56, 31] exploited this by incorporating depth maps into a tracking framework to improve robustness in crowded scenes. Similarly, thermal infrared imaging captures thermodynamic signatures independent of visible light, proves effective in low-illumination. [62, 19, 27, 52] demonstrated that fusing thermal and RGB data can yield more reliable appearance representations.
@@
In summary, contemporary methods emphasized effective multi-modal interaction and fusion [42, 45, 59, 65]. With the emergence of large-scale RGB datasets and universal backbones (e.g., vision transformer), pre-trained trackers [58, 49, 9] have demonstrated remarkable generalization across diverse scenarios. These advances have shifted the paradigm toward transferring knowledge from pre-trained models in designing high-performance trackers. Consequently, multi-modal tracking is increasingly driven by reusing RGB-induced semantic priors, particularly when annotated multi-modal data are scarce or noisy. In this work, we target this imperative by optimizing the adaptation of pre-trained trackers for efficacious cross-modality transfer.
II-B Cross-modal Transfer Learning
@@
To adapt pre-trained models for multi-modal tracking, two main transfer learning strategies have recently emerged. Some works adhere to the full fine-tuning (FFT) paradigm [42, 45, 73, 41], where pre-trained models serve as weight initializations and are entirely re-trained on the target tasks. These methods require a shared or compact cross-modal feature space to inherit the generalization capability of the original model. Representatively, SUTrack [9] employed a single vision transformer with unified input representations for multiple modalities, avoiding task-specific customization. While effective, one primary dilemma may be innate to FFT: the contradiction between the paucity of large-scale downstream datasets and the huge appetites of cross-domain adaptation. This imbalance frequently leads to catastrophic forgetting of the pre-trained knowledge and severe overfitting to the limited target data. To alleviate this, profiting from the affluent experience of natural language processing and computer vision communities [21, 7, 18], the field has pivoted toward parameter-efficient fine-tuning (PEFT). The core principle of PEFT is to keep the majority of pre-trained weights frozen while tuning only a minimal number of additional parameters dedicated to the new task. By lightly fine-tuning, these methods aim to preserve the generalization while mitigating overfitting. For example, ProTrack [57] and ViPT [70] modulated RGB features by introducing trainable auxiliary tokens into attention layers. Similarly, methods like BAT [5] and SDSTrack [17] inserted lightweight adapter modules between attention layers for cross-modal shift compensation. More recently, unified trackers like UnTrack [51] and OneTracker [16] employed LoRA and prompt tuning to seamlessly integrate multiple modalities, enabling effective unification across diverse inputs. Despite their effectiveness, PEFT methods suffer from intrinsic limitations. First, the rigid constraints can lead to underfitting, limiting the model’s capacity to handle vast distribution drifts. Conversely, the randomly initialized add-on modules lack alignment with the pre-trained manifold, creating an optimization gap that predisposes the model to overfitting. Consequently, achieving robust and reliable fine-tuning for multi-modal tracking remains an elusive goal.
@@
Remark. While recent works have explored sensitivity-aware mechanisms in efficient fine-tuning (SPT) and continual learning (CL), our SRFT introduces a fundamentally distinct paradigm in objectives and mechanisms. Specifically, methods like SPT [15] and CDRA-SPT [8] utilized sensitivity to prioritize structural adaptation, such as parameter masking or dynamic rank allocation for parameter efficiency. Conversely, MACL [47] employed sensitivity to prevent forgetting in sequential tasks by minimizing loss variance. Crucially, these approaches typically operate with an exclusive focus on the target or current data. Diverging from this, SRFT establishes a gradient-regularized paradigm. We formulate sensitivity to balance the preservation of pre-trained knowledge with sparse transfer gradient, via a hybrid sensitivity from pre-trained to multi-modal contexts. This distinction renders our method orthogonal to existing FFT, PEFT, and SPT frameworks.
@@
Figure 2: Network architecture of the multi-modality trackers. All modules are initialized with the weights of a pre-trained RGB tracker. We only fine-tune backbones and fusion blocks with sensitivity-aware regularization terms.
III Proposed Method
@@
Learning generalized and coherent representations is crucial for adapting RGB-based models to multi-modal trackers. To unlock the full potential of pre-trained trackers, we revisit the core principles governing the cross-modal transfer process. In Section III-A, we first present the architectural design of our multi-modal trackers and highlight the key challenges associated with full fine-tuning. Next, Section III-B identifies and quantifies pronounced prior and transfer sensitivities that characterize pre-trained knowledge and cross-domain adaptation. Finally, in Section III-C, we incorporate these sensitivities into the optimization phase, enabling adaptive and dynamic modulation of parameter updates.
III-A Preliminaries
@@
Network Architecture of Multi-modality Trackers. Fig. 2 depicts the overall architecture of our multi-modal tracker, in which all modules are initialized with the weights of a pre-trained RGB tracker. For a multi-modal tracker, the RGB and auxiliary inputs are first fed to the embedding layer to generate the corresponding tokens. Then, symmetric transformer backbones (e.g., ViT or its variants) handle feature extraction and interaction. Without involving customized multi-modal fusion modules, we repurpose certain ViT blocks for multi-stage fusion of multi-modal tokens. Finally, the fused features are fed into the box head to estimate the object state. To retain modal-agnostic object association knowledge, the pre-trained box head is utilized and kept frozen.
@@
Learning Process of Cross-Domain Transfer. We consider $f_{\theta}(\cdot):X\rightarrow Y$ as a multi-modal tracker with parameters $\theta\in\mathbb{R}^{|\theta|}$, with $|\theta|$ representing the total number of the parameters $\theta$. It takes task dataset $D=\{X,Y\}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ to yield the optimal $\theta$, where $x$ represents a multi-modal input pair and $y$ is the target bounding box. In this paper, we focus on efficiently transferring the models pre-trained on a source domain $D_{0}$ (e.g., large-scale RGB data) to downstream domains $D_{t}$ (e.g., auxiliary modalities or degraded images), where typically $\left|D_{0}\right|\gg\left|D_{t}\right|$. In this setting, let $\theta_{0}$ be the pre-trained parameters (optimal on $D_{0}$), which serve as initialization. After fine-tuning on $D_{t}$, the parameters become $\theta_{t}$. Given a task-specific loss function $\mathcal{L}$, the vanilla transfer objective is:
@@
However, directly optimizing $\theta_{t}$ as per Eq. (1) leads to severe overfitting and unstable tuning, as evidenced by the pronounced train-test performance gap and oscillations in training dynamics observed in Fig. 1. This motivates a more principled transfer learning approach, as described next.
@@
III-B Modeling Sensitivity for Multi-Modality Tracking
Effective cross-domain fine-tuning necessitates a delicate balance between stability (preserving critical pre-trained knowledge) and plasticity (adapting to the new domain) [34]. We regulate this trade-off via two intrinsic parameter sensitivities, identifying the evolving parameter importance from pre-trained to target tasks: prior sensitivity, reflecting its importance to pre-trained knowledge, and transfer sensitivity, indicating its role in target-task adaptation. Excessive changes to highly sensitive parameters in either aspect can hinder the transfer process.
Manifold hypothesis. As shown in Fig. 3, the parameter set $\{\theta\}_{L=L_{c}}$, with certain loss $L_{c}$, does not fill the entire space $\mathbb{R}^{|\theta|}$. Instead, it usually lies in a manifold $\mathcal{M}_{L_{c}}$, which is also continuous in $\mathbb{R}^{|\theta|}$ [13, 40].
Based on the aforementioned manifold hypothesis, to avoid the loss of generalization ability for fine-tuned models on downstream domains, the critical thing lies in optimizing the model in local tangent space, shown as the red arrow of Fig. 3. Thus, in the following section, we investigate the method to approximate and optimize the model on such tangent space.
III-B1 Safeguarding Generalization via Prior Sensitivity
In this context, we first analyze the influence of the tangent space on model generalization from the pre-trained loss-parameter manifold, and then design an efficient eigen-decomposition method to quantify this sensitivity.
Exploring Parameter Manifold Geometry. We first define the joint empirical risks for tuning models over the union of pre-training and downstream tasks as

### v3

@@ Abstract
This paper tackles the critical challenge of optimizing multi-modality trackers by effectively adapting pre-trained models for RGB data. Existing fine-tuning paradigms oscillate between excessive flexibility and over-restriction, both leading to suboptimal plasticity-stability trade-offs. To mitigate this dilemma, we propose a novel significance-regularized fine-tuning framework, which delicately refines the learning process by incorporating intrinsic parameter significance. Through a comprehensive investigation of the transition from pre-trained to multi-modality contexts, we identify that parameters crucial to preserving foundational patterns and managing cross-domain shifts are the primary drivers of this issue. Specifically, we first probe the tangent space of pre-trained weights to measure and orient prior significance, dedicated to preserving generalization. Subsequently, we characterize transfer significance during the fine-tuning phase, emphasizing adaptability and stability. By incorporating these parameter significance terms as unified regularization, our method markedly enhances transferability across modalities. Extensive experiments showcase the superior performance of our method, surpassing current state-of-the-art techniques across various multi-modal tracking benchmarks. The source code and models are publicly available at https://github.com/zhiwen-xdu/SRTrack.
journal: International Journal of Computer Vision
@@ 1 Introduction
Object tracking, a foundation task of visual perception, has witnessed remarkable advancements over the past decades (Hong et al., 2024; Xie et al., 2024; Zheng et al., 2024a; Cai et al., 2024). Despite the promising results, RGB-based trackers often struggle with some complex and degraded conditions, such as extreme illumination, motion blur, and occlusions. Therefore, multi-modality tracking with more comprehensive sensory signals (e.g., event, depth, thermal) has garnered growing interest. With the popularity of the data-driven methods in the object tracking community, both data scale and model capacity have experienced huge explosions in recent years (Ye et al., 2022; Lin et al., 2022; Chen et al., 2023, 2025b). There is a prevailing paradigm that explores these pre-trained trackers on large-scale RGB-based datasets and adapts them to diverse auxiliary modalities, a process known as cross-modal fine-tuning or transfer learning, to enhance performance and accelerate convergence by starting from well-structured pre-trained weights.
@@
Concretely, some existing approaches have explored the full fine-tuning (FFT) paradigm (Wang et al., 2023; Zhu et al., 2023c; Sun et al., 2025), where multi-modality trackers are initialized with RGB-based weights and subsequently optimized for task-specific objectives. Nevertheless, while FFT offers maximum flexibility and swiftly adapts to the target domain, substantial domain gaps coupled with limited scale of auxiliary modalities, hinder the retention of pre-trained knowledge structure during transfer, often inducing severe overfitting (i.e., a widened train-test gap and deteriorating test performance). In contrast to full fine-tuning, recent research has shifted toward parameter efficient fine-tuning (PEFT) (Hu et al., 2021; Jia et al., 2022; Chen et al., 2022). The PEFT keeps the majority of pre-trained parameters frozen, updating only a small fraction of modality-specific ones to retain prior knowledge. Several methods fall under this umbrella (Zhu et al., 2023a; Hou et al., 2024; Wu et al., 2024b), including prompt tuning, visual adapter, etc. Although effective, PEFT-based methods impose rigid constraints on the primary model weights, resulting in underfitting (i.e., a restricted upper bound on training performance) when handling the significant distribution drifts. In summary, existing fine-tuning paradigms fluctuate between excessive flexibility and over-restriction, both contributing to misfitting (i.e., overfitting and underfitting) and a sub-optimal plasticity-stability trade-off between pre-trained knowledge and downstream adaptation (as clearly illustrated in Fig. 1).
@@
In this work, we endeavor to mitigate the misfitting dilemma in cross-modal tracker adaptation by strategically regularizing the learning process. By analyzing parameter responses spanning from pre-training to fine-tuning phases, we identify that pronounced parameter significance, in its various aspects, is the primary cause of degraded prior generalizability and transfer adaptability. To mitigate this, we propose a significance-regularized fine-tuning (SRFT) framework that delicately calibrates the gradient updates to boost model transfer with precision. Specifically, we optimize multi-modality trackers from the following perspectives.
  1. 1.
@@
    Formulating Prior Significance. We commence by investigating the tangent space of pre-trained parameters, utilizing it as a critical indicator to assess and preserve prior generalizability for downstream fine-tuning. Furthermore, we estimate this significance via an eigen-decomposition approximation.
  2. 2.
@@
    Modeling Transfer Significance. To further elucidate the adaptation challenges, we explore how sparse gradients exacerbate fine-tuning instability. Accordingly, we formulate transfer significance using off-the-shelf gradient matrices, aiming to facilitate effective gradient rebalancing.
  3. 3.
@@
    Significance Regularized Tuning. By harnessing the identified parameter significance, we suggest an adaptive tuning scheme that safeguards essential pre-trained knowledge and fosters coherent multi-modality representations through finely modulated, significance-driven updates. This mechanism facilitates seamless fine-tuning across various multi-modality tracking tasks, continuously enhancing the model during the training phase.
@@
Our method strategically guides the cross-domain fine-tuning process to optimize downstream multi-modality tracking tasks. Extensive experimental results showcase our method achieves new state-of-the-art results across three multi-modality tracking tasks (RGB-Event, RGB-Depth, RGB-Thermal) and seven benchmarks, spanning diverse pre-trained tracking models. Comprehensive ablation studies and parameter significance measurements confirm the effectiveness of the significance-aware regularization fine-tuning strategy. In summary, the main contributions of this paper are:
@@
  • we revisit the misfitting issue of multi-modality tracking for adapting foundation models and propose a novel regularized tuning framework (SRFT) to indicate better transferability, which is orthogonal to the existing FFT and PEFT methods;
@@
  • we formulate the parameter significance with respect to pre-trained knowledge and transfer stability, and introduce a significance-aware update strategy to refine the learning process, thereby facilitating the generalization and adaptability of trackers;
@@
The remainder of the paper is organized as follows. Section 2 reviews existing literature on multi-modal trackers and cross-modal transfer learning methods. In Section 3, we rigorously formulate prior and transfer parameter significance from both pre-training and fine-tuning perspectives, followed by the implementation of significance-regularized tuning. Section 4 presents extensive experiments to demonstrate the effectiveness of our method, along with comprehensive ablation studies to analyze the impact of different components and designs. Finally, Section 5 concludes the paper.
2 Related Work
2.1 Multi-Modal Object Tracking
@@
Object tracking involves localizing an object across frames given its initial appearance (Wei et al., 2023; Bai et al., 2024; Lin et al., 2024; Zheng et al., 2024b). However, the vulnerability of RGB-only trackers under adverse conditions has driven the development of multi-modal tracking, where auxiliary cues complement the intrinsic deficiencies of visible imagery (Zhang et al., 2023c). For instance, event cameras provide robust dynamic information under extreme motion or lighting, enabling RGB-event fusion for high-speed and low-dynamic tracking (Zhu et al., 2022; Chen et al., 2024b; Wang et al., 2024; Zhang et al., 2024b; Wang et al., 2025a, b). Depth offers geometric priors that help handle occlusion and background clutter, and has been integrated to improve robustness in crowded scenes (Lukezic et al., 2019; Qian et al., 2021; Yan et al., 2021b; Liu et al., 2025). Thermal infrared imaging captures thermodynamic signatures independent of visible light, proves effective in low-illumination. (Zhang et al., 2021b; Hui et al., 2023; Li et al., 2025; Xiang et al., 2025) demonstrated that fusing thermal and RGB data can yield more reliable appearance representations.
@@
In summary, contemporary methods emphasized effective multi-modal interaction and fusion (Zhang et al., 2024a, b; Chen et al., 2024a; Tan et al., 2025a, b; Feng et al., 2025). With the emergence of large-scale RGB datasets and universal backbones (e.g., vision transformer), pre-trained trackers (Ye et al., 2022; Wu et al., 2023; Chen et al., 2025b) have demonstrated remarkable generalization across diverse scenarios. These advances have shifted the paradigm toward transferring pre-trained models for designing high-performance multi-modal trackers. Consequently, multi-modal tracking is increasingly driven by reusing RGB-induced semantic priors, particularly when annotated multi-modal data are scarce or noisy. In this work, we target this imperative by optimizing the adaptation of pre-trained trackers for efficacious cross-modal transfer learning.
2.2 Cross-modal Transfer Learning
@@
To adapt pre-trained models for multi-modal tracking, two main transfer learning strategies have recently emerged. Some works adhere to the full fine-tuning (FFT) paradigm (Tang et al., 2022; Wang et al., 2023; Zhu et al., 2023c; Sun et al., 2025), where pre-trained models serve as weight initializations and are entirely re-trained on the target tasks. These methods require a shared or compact cross-modal feature space to inherit the generalization capability of the original model. Representatively, SUTrack (Chen et al., 2025b) employed a single vision transformer with unified input representations for multiple modalities, avoiding task-specific customization. While effective, one primary dilemma may be innate to FFT: the contradiction between the paucity of large-scale downstream datasets and the huge appetites of cross-domain adaptation. This mismatch frequently leads to catastrophic forgetting of the pre-trained knowledge and severe overfitting to the limited target data.
To alleviate this, profiting from the affluent experience of natural language processing and computer vision communities (Jia et al., 2022; Chen et al., 2022; Hu et al., 2021), the field has pivoted toward parameter-efficient fine-tuning (PEFT). The core principle of PEFT is to keep the majority of pre-trained weights frozen while tuning only a minimal number of additional parameters dedicated to the new task. By lightly fine-tuning, these methods aim to preserve the generalization while mitigating overfitting. For example, ProTrack (Yang et al., 2022) and ViPT (Zhu et al., 2023a) modulated RGB features by introducing trainable auxiliary tokens into attention layers. Similarly, methods like BAT (Cao et al., 2024) and SDSTrack (Hou et al., 2024) inserted lightweight adapter modules between attention layers for cross-modal shift compensation. More recently, unified trackers like UnTrack (Wu et al., 2024b) and OneTracker (Hong et al., 2024) employed LoRA and prompt tuning to seamlessly integrate multiple modalities, enabling effective unification across diverse inputs. Despite their effectiveness, PEFT methods suffer from intrinsic limitations. First, the rigid constraints can lead to underfitting, limiting the model’s capacity to handle vast distribution drifts. Conversely, the randomly initialized add-on modules lack alignment with the pre-trained manifold, creating an optimization gap that predisposes the model to overfitting. Consequently, achieving robust and reliable fine-tuning for multi-modal tracking remains an elusive goal.
@@
Remark. Beyond fine-tuning strategies in multi-modal tracking, we also review related fine-tuning paradigms from adjacent domains that could be conflated with ours, to more clearly distinguish our contribution. In particular, some recent studies have explored parameter sensitivity in efficient fine-tuning (SPT) and continual learning (CL). Notably, our method differs fundamentally from these approaches in both objective and mechanism. Specifically, SPT (He et al., 2023) and CDRA-SPT (Chen et al., 2025a) leveraged parameter sensitivity to guide structural adaptation for efficiency, e.g., sparse parameter updates or dynamic rank allocation. MACL (Wang and Huang, 2024) used sensitivity to mitigate forgetting in sequential tasks by minimizing loss variance. Crucially, these approaches typically operate with an exclusive focus on the target context. Diverging from this, SRFT adopts a gradient-regularized paradigm: we define a hybrid parameter significance spanning pre-trained and cross-modal domains to balance prior generalization with target adaptability. This distinction renders our method orthogonal to existing FFT, PEFT, and SPT methods.
@@
Figure 2: Network architecture of our multi-modality trackers. All modules are initialized with the weights of a pre-trained RGB tracker. We only fine-tune backbones and fusion blocks with significance-aware regularization terms.
3 Proposed Method
@@
Learning generalized and coherent representations is crucial for adapting RGB-based models to multi-modal trackers. To unlock the full potential of pre-trained trackers, we revisit the core principles governing the cross-modal transfer process. In Section 3.1, we first present the architectural design of our multi-modal trackers and highlight the key challenges associated with full fine-tuning. Next, Section 3.2 identifies and quantifies pronounced prior and transfer parameter significance that characterize pre-trained knowledge and cross-domain adaptation. Finally, in Section 3.3, we incorporate these significance measurements into the optimization phase, enabling adaptive and dynamic modulation of parameter updates.
3.1 Preliminaries
@@
Network Architecture of Multi-modality Trackers. Fig. 2 depicts the architecture of our multi-modal trackers, in which all modules are initialized with the weights of a pre-trained RGB tracker. For the multi-modal tracker, RGB and auxiliary inputs are first fed to the embedding layer to generate the corresponding template and search tokens. Then, symmetric transformer backbones (e.g., ViT or its variants) handle feature extraction and interaction. Without involving customized multi-modal fusion modules, we repurpose certain ViT blocks (e.g., layers 2, 5, 8, and 11) for multi-stage fusion by concatenating multi-modal template and search tokens. Finally, the fused features are fed into the box head to estimate the object state. To retain modal-agnostic object association knowledge, the pre-trained box head is utilized and kept frozen. Additional model details can be found in Appendix A.
@@
Learning Process of Cross-Domain Transfer. We consider $f_{\theta}(\cdot):X\rightarrow Y$ as a multi-modal tracker with parameters $\theta\in\mathbb{R}^{|\theta|}$, with $|\theta|$ representing the total number of the parameters $\theta$. The model takes a task-specific dataset $D=\{X,Y\}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ to yield the optimal $\theta$, where $x_{i}$ represents a multi-modal input pair and $y_{i}$ is the corresponding object bounding box or output for the task. In this work, we focus on efficiently transferring the models that have been pre-trained on a source domain $D_{0}$ (e.g., large-scale RGB data) to downstream domains $D_{t}$ (e.g., auxiliary modalities or degraded versions of images), where typically $\left|D_{0}\right|\gg\left|D_{t}\right|$. Let $\theta_{0}$ represent the pre-trained parameters, which are optimal for the source domain $D_{0}$, and these parameters are used as the weight initialization. After fine-tuning the model on the target domain $D_{t}$, the model parameters become $\theta_{t}$. Given a task-specific loss function $\mathcal{L}$, the vanilla transfer objective is formulated as follows:
@@
However, directly optimizing $\theta_{t}$ as per Eq. (1) leads to severe overfitting and unstable tuning, as evidenced by the pronounced train-test performance gap and oscillations in training dynamics observed in Fig. 1. This occurs because the model tends to over-adapt to the unstable target domain $D_{t}$, while neglecting the generalization capabilities learned from the source domain $D_{0}$. These challenges motivate a more principled transfer learning approach, as described next.
@@
3.2 Modeling Significance for Multi-Modality Tracking
Effective cross-domain fine-tuning necessitates a delicate balance between stability (preserving critical pre-trained knowledge) and plasticity (adapting to the new domain) (Mermillod et al., 2013; Zheng et al., 2025; Zhou et al., 2025). We regulate this trade-off via two intrinsic parameter significance, identifying the evolving parameter importance from pre-trained to target tasks: prior significance, reflecting its importance to pre-trained knowledge, and transfer significance, indicating its role in target-task adaptation. Excessive changes to highly significant/fragile parameters in either aspect can hinder the transfer process.
Loss-Parameter Manifold Hypothesis. As shown in Fig. 3, the parameter model set $\{\theta\}_{L=L_{c}}$, with certain loss $L_{c}$, does not fill the entire space $\mathbb{R}^{|\theta|}$. Instead, it usually lies in a manifold $\mathcal{M}_{L_{c}}$, which is also continuous in $\mathbb{R}^{|\theta|}$ Bengio et al. (2013); Fefferman et al. (2016); Song and Ermon (2019); Meilă and Zhang (2024). Based on the manifold hypothesis, to avoid the loss of generalization ability for fine-tuned models on downstream domains, the critical thing lies in optimizing the model in local tangent space, shown as the red arrow of Fig. 3. Thus, in the following section, we investigate the method to approximate and optimize the model on such tangent space.
3.2.1 Safeguarding Generalization via Prior Significance
In this context, we first analyze the influence of the tangent space on model generalization from the pre-trained loss-parameter manifold, and then design an efficient eigen-decomposition method to quantify this prior significance.
Exploring Loss-Parameter Manifold Geometry. We first define the joint empirical risks for tuning models over the union of pre-training and downstream tasks as: